OpenStudy (anonymous):

dx/dt=20sec^2(theta) dtheta/dt When isolating dtheta/dt, the sec^2(theta) turns into cos^2(theta). Can someone refresh my drink as to why this is so?

OpenStudy (psymon):

Well, you divide both sides by sec^2(theta) and youd get \[\frac{ dx }{ \sec^{2} \theta dt }\]But by identity, 1/secx = cosx, so it can simply be rewritten as \[\frac{ dxcos^{2} \theta }{ dt }\]

OpenStudy (anonymous):

gentleman and a scholar

OpenStudy (dan815):